OpenStudy (anonymous):

A box has a width of 1 inch less than the length and a height 4 inches greater than the length. The volume of the box is 12 in cubed. Find the length of the box

OpenStudy (ddcamp):

V = l*w*h w = l-1 h = l+4 V = l*(l-1)*(l+4)

OpenStudy (anonymous):

So where is the length found?

OpenStudy (ddcamp):

Factor the equation to get:  The problem says that V=12